Three sisters, Mimi, Alison and Tracy decide to buy a house. Mimi pays

1

_

3

of the cost, Alisa pays

2

_

5

of

the cost and Tracy pays the rest.



2.2.1 What fraction of the cost of the house does Tracy pay? [2]
1
Expert's answer
2020-06-02T18:12:36-0400

In order to find the fraction of Tracy, we have to subtract 1/3 and 2/5 from 1.

As we know, 1 equals to 15/15. In this way, 1/3 equals to 5/15 and 2/5 equals to 6/15.

Thus, 15/15 - 5/15 - 6 /15 = (15 - 5 - 6) / 15 = 4 / 15

The answer is 4 / 15